# Recurrent Self-Attention Dynamics: An Energy-Agnostic Perspective from Jacobians

### v2

@@ B.1 Experimental setup
In our experiments, we used two Sudoku datasets: the SATNet [Wang et al., 2019] and RRN dataset [Palm et al., 2018]. The key differences between the two are that the RRN dataset is more difficult (with only 1717171734343434 given digits compared to 3131313142424242 in SATNet) and larger in size ($198$k samples vs. $10$k samples). Following Miyato et al. [2025], we used the SATNet dataset for training as in-distribution (ID) data and the RRN dataset as out-of-distribution (OOD) data. This setup allows us to evaluate the ability of models to generalize to more challenging settings.
We primarily followed Miyato et al. [2025] and used their official implementation [^1] https://github.com/autonomousvision/akorn, using $N=4$ as the dimension of oscillators. We used the Adam optimizer [Kingma and Ba, 2015] and trained for $100$ epochs with batch size $100$. We tuned the learning rate across $\{1\times 10^{-6},5\times 10^{-6},\dotsc,1\times 10^{-3}\}$ for all settings based on the OOD accuracy at the iteration $T=16$. All experiments were conducted on NVIDIA H200 GPUs, and we run experiments with 5 different random seeds.
We also conducted experiments on the CIFAR-10 dataset [Krizhevsky et al., ]. See table S.1 for training and model configurations.
@@ B.2 Single-head generalized symmetric SA
under the condition that $H=1$. If $R_{\text{E-single}}=0$, setting $\bm{W}^{V}=\bm{W}_{1}^{V}\bm{W}_{1}^{O}$ satisfies the condition on $\bm{W}^{V}$ described in Proposition 4.1.
@@ B.4 Details of other figures
For Figure 2, we computed the eigenvalues of the Jacobian matrix at $T=16$ on a randomly selected sample from the Sudoku dataset. For the model with normalization, we used the fully trained model. For models without normalization, we followed the same setup as in the Lyapunov experiments and used models trained for only one epoch.
For the computation of the SA’s Jacobian in Figure 1, we used the CCDV arXiv summarization dataset [Cohan et al., 2018], as it provides text data suitable for varying the number of tokens. We used an initialized SA and computed the Jacobian and SA followed by normalization over $500$ randomly selected samples. The norm of tokens was set to $R=100$ and their dimensions to $D=256$.
@@ B.5 Additional results
(a) Number of attention heads $H$ vs. Lyapunov exponent in ItrSA with normalization.
(b) Number of attention heads $H$ and OOD accuracy on the Sudoku dataset.
Figure S.1: Effect of the number of attention heads $H$ on Lyapunov exponent and OOD accuracy in the Sudoku dataset.
The number of attention heads. To further investigate the effect of normalization, we calculated the Lyapunov exponents while varying the number of attention heads. The results in Figure S.1 indicate that the number of heads has little to no impact on the Lyapunov exponents. Figure S.1 also shows the OOD test accuracy for models with different numbers of attention heads. The results indicate that models with a small number of heads ($H=1,2$) exhibit poor performance, while models with $H=8$ or $16$ achieve the highest accuracy.
@@
Table S.2 presents the values of the $\bm{\gamma}$ parameter learned by ItrSA. The results indicate that the trained models exhibit small $\bm{\gamma}$ values, with $\max_{j}|\gamma_{j}|<1$.
Table S.2: $\bm{\gamma}$ in $\operatorname{RMSNorm}$ with different $N$.
$N$ $\|\bm{\gamma}\|$ $\max_{j}|\gamma_{j}|$
$4$ $0.229\pm 0.0004$ $0.229\pm 0.0004$
$8$ $0.031\pm 0.0017$ $0.052\pm 0.0000$
$16$ $0.098\pm 0.0062$ $0.098\pm 0.0062$
$32$ $0.348\pm 0.0004$ $0.348\pm 0.0004$
$64$ $0.489\pm 0.0006$ $0.489\pm 0.0006$
$128$ $0.841\pm 0.0000$ $0.841\pm 0.0000$
$256$ $0.738\pm 0.0004$ $0.738\pm 0.0004$
$512$ $0.811\pm 0.0002$ $0.811\pm 0.0002$
Further results on the interpretation of pseudo-energy via the Jacobian
We provide additional insights into the interpretation of pseudo-energy via the Jacobian discussed in Section 6.3. Using Lemma A.3 from Noci et al. [2022], the Jacobian of SA is expressed as
@@
Figure S.2 empirically confirmed that both the contribution index and the pseudo-energy behave similarly even under these approximations. PV indicates the contribution index using the approximation (93), and PV0 uses (94). The figure shows that the contribution index can be effectively explained using only the attention matrix at the initialization of inference. The blue curve shows the approximated pseudo-energy (95) and works similarly to the original one. Thus, our Jacobian-based analysis interprets the pseudo-energy as quantifying exploration in eigenspaces corresponding to large eigenvalues of the attention matrix determined by the initial inference states.
@@
B.6 CIFAR-10
For the experiments on the CIFAR-10 dataset [Krizhevsky et al., ], we used the same architecture and setup as in the Sudoku experiments. We used the Adam optimizer and tuned the learning rate across $\{1\times 10^{-6},5\times 10^{-6},\dotsc,1\times 10^{-3}\}$ based on the test accuracy at the iteration $T=16$. We trained models for 200 epochs and set the batch size $128$.
Figure S.3 shows the Lyapunov exponent on the CIFAR-10 dataset. This result is in the same trend with that on the Sudoku dataset.
@@ C.1 Extended related work
The Transformer architecture has been a focus of efforts to provide theoretical grounding. Geshkovski et al. [2023a, b, 2024] formulated recurrent SA dynamics as interactions among tokens (“particles”), enabling theoretical analysis of phenomena such as meta-stable clustering and rank-one collapse. Their continuous-time dynamics monotonically decrease an energy (Lyapunov) function given by a summation over exponential functions, commonly requiring constraints such as single-head attention, hyperspherical token states, and symmetric weights. Karagodin et al. [2024] extended this framework to the case of causal attention masking. Bruno et al. [2025] succeeded in mathematically characterizing the meta-stable clustering as a Wasserstein gradient flow of mean-field token dynamics, with the energy serving as its potential function, although they replaced the softmax function with an unnormalized exponential function and restricted their analysis to identity weight matrices. Yang et al. [2022] considered an exponential energy function similar to that of Geshkovski et al. [2023a], describing the Transformer as performing alternating majorization-minimization updates on distinct energy functions. Their approach also accommodates discrete state updates and MLP layers, although it entails complex conditions, including constraints on step sizes and proximity to fixed points. Ramsauer et al. [2021] formalized the cross-attention mechanism as modern Hopfield networks, Hoover et al. [2023], Hu et al. [2025] further developed energy functions for Transformers including self-attentions. We do not address approaches based on Hopfield networks in this work, as they require architectural modifications, such as adding auxiliary signal paths that are absent in standard Transformers, which are beyond our scope.

### v3

@@ B.1 Experimental setup
In our experiments, we used two Sudoku datasets: the SATNet [Wang et al., 2019] and RRN dataset [Palm et al., 2018]. The key differences between the two are that the RRN dataset is more difficult (with only 17173434 given digits compared to 31314242 in SATNet) and larger in size (198198k samples vs. 1010k samples). Following Miyato et al. [2025], we used the SATNet dataset for training as in-distribution (ID) data and the RRN dataset as out-of-distribution (OOD) data. This setup allows us to evaluate the ability of models to generalize to more challenging settings.
We primarily followed Miyato et al. [2025] and used their official implementation [^1] https://github.com/autonomousvision/akorn, setting the dimension of oscillators of AKOrN to $N=8$. The readout module of AKOrN (described in Section˜C.2) was also incorporated into our ItrSA model. We used the Adam optimizer [Kingma and Ba, 2015] and trained for $100$ epochs with batch size $100$. For all settings, we tuned the learning rate over $\{1\times 10^{-6},5\times 10^{-6},\dotsc,1\times 10^{-3}\}$ and, for regularization methods in Figure˜5, the parameter $\lambda$ over $\{1\times 10^{-8},1\times 10^{-7},\dotsc,1\times 10^{-1}\}$, selecting values based on OOD accuracy at iteration $T=16$. All experiments were conducted on NVIDIA H200 GPUs, and we run experiments with 5 different random seeds.
We also conducted experiments on the CIFAR-10 dataset [Krizhevsky et al., ]. See table˜S.1 for training and model configurations.
@@
Table S.2: Regularization coefficients used in Figure˜5.
Method Value
E-single 1e-6
E-multi 1e-4
Spec (ItrSA) 1e-4
Spec (AKOrN) 1e-5
@@ B.2 Single-head generalized symmetric SA
under the condition that $H=1$. If $R_{\text{E-single}}=0$, setting $\bm{W}^{V}=\bm{W}_{1}^{V}\bm{W}_{1}^{O}$ satisfies the condition on $\bm{W}^{V}$ described in Proposition˜4.1.
@@ B.3 Lyapunov exponent
To mitigate numerical sensitivity, we also use mean Lyapunov exponent
$\lambda_{\text{mean}}\coloneqq\frac{1}{M}\sum_{i=1}^{M}\lambda_{i},$ (115)
where $M$ denotes the number of Lyapunov exponents.
@@ B.4 Details of other figures
For Figure˜2, we computed the eigenvalues of the Jacobian matrix at $T=16$ on a randomly selected sample from the Sudoku dataset. For the model with normalization, we used the fully trained model. For models without normalization, we followed the same setup as in the Lyapunov experiments and used models trained for only one epoch.
For the computation of the SA’s Jacobian in Figure˜1, we used the CCDV arXiv summarization dataset [Cohan et al., 2018], as it provides text data suitable for varying the number of tokens. We used an initialized SA and computed the Jacobian and SA followed by normalization over $500$ randomly selected samples. The norm of tokens was set to $R=100$ and their dimensions to $D=256$.
@@ B.5 Additional results
B.5.1 Effects of model structure and regularization
Number of attention heads. To further examine the effect of normalization, we computed Lyapunov exponents while varying the number of attention heads $H$. Because the maximum Lyapunov exponent is numerically sensitive, we report the mean Lyapunov exponent instead. The results in Figure˜S.1 show that models with few heads ($H\in\{1,2\}$) perform poorly, whereas $H=8$ achieve the highest accuracy. Moreover, the mean Lyapunov exponent is positively correlated with accuracy, suggesting that more dynamic states are associated with better performance.
(a) ID
(b) OOD
Figure S.1: Effect of the number of attention heads $H$ on mean Lyapunov exponent and accuracy in the Sudoku dataset.
@@ B.5.1 Effects of model structure and regularization
Table˜S.3 presents the values of the $\bm{\gamma}$ parameter learned by ItrSA. The results indicate that the trained models exhibit small $\bm{\gamma}$ values, with $\max_{j}|\gamma_{j}|<1$.
Table S.3: $\bm{\gamma}$ in $\operatorname{RMSNorm}$ with different NN.
NN $\|\bm{\gamma}\|$ $\max_{j}|\gamma_{j}|$
44 $0.229\pm 0.000$ $0.229\pm 0.000$
88 $031\pm 0.002 0.0$$052\pm 0.000$
$16$ $098\pm 0.006 0.0$$098\pm 0.006$
$32$ $0.348\pm 0.000$ $0.348\pm 0.000$
$64$ $0.489\pm 0.001$ $0.489\pm 0.001$
$128$ $0.841\pm 0.000$ $0.841\pm 0.000$
$256$ $0.738\pm 0.000$ $0.738\pm 0.000$
$512$ $0.811\pm 0.000$ $0.811\pm 0.000$
Further results on regularization.
In Figure˜S.2, we evaluate the effect of the proposed multi-head energy (Proposition˜4.2) on both the mean Lyapunov exponent and accuracy. As the regularization strength $\lambda$ increases, both metrics consistently decrease. Even at $\lambda=0$, accuracy is lower than the original model due to the orthogonality constraint. The “Hard” constraint yields the lowest accuracy and the smallest Lyapunov exponent. Overall, these results indicate that stronger regularization suppresses the Lyapunov exponent but also degrades accuracy, suggesting that while multi-head energy regularization encourages more convergent dynamics, it does not necessarily yield better performance.
(a) ID
(b) OOD
Figure S.2: Effect of multi-head energy on the mean Lyapunov exponent and accuracy for the Sudoku dataset. $\lambda$ denotes the regularization coefficient of Eq. (20), and “Hard” indicates the hard constraint in Proposition˜4.2.
In Figure˜S.3 we plot the case where we used $N=8$ as the oscillator dimension of AKOrN. AKOrN achieves the best performance and spectral regularization is also effective.
Figure S.3: Energy-based regularization (“E-single” and “E-multi”) underperforms the original methods, while Jacobian spectral regularization (“Spec”) outperforms. We used $N=8$ for AKOrN.
B.5.2 Distribution of Lyapunov exponents
Figure˜S.4 shows the distibution of the Lyapunov exponent.
(a) AKOrN
(b) ItrSA
Figure S.4: Distribution of the Lyapunov exponent.
B.5.3 Lyapunov exponents in language modeling tasks
To evaluate in a more realistic scenario, we conducted language modeling experiments on the BabyLM Challenge dataset (2023, 10M version) [Diehl Martinez et al., 2023], comparing our ItrSA model with GPT-2 [Radford et al., 2019]. We trained for 30 epochs using the AdamW optimizer. As shown in Table˜S.4, the maximum Lyapunov exponent (MLE) for ItrSA is slightly positive, which is consistent with our other tasks and suggests mildly chaotic dynamics. We also show the loss values in the case of GPT-2 as a reference which confirms that the performance of our ItrSA can become comparable to that of GPT-2.
Table S.4: Language modeling results. MLE denotes the maximum Lyapunov exponent.
Model # Parameters Training loss Validation loss MLE
ItrSA 96.2M 4.31 5.67 $0.196\pm 0.016$
GPT-2 124M 1.49 5.58 -
B.5.4 Sensitivity to initial conditions
In our study, we define “criticality” as the point at which the largest Lyapunov exponent takes zero. In our experiments, when the maximum Lyapunov exponent gets slightly positive, we observed behaviors consistent with widely accepted notions of chaos, particularly sensitivity to initial conditions as follows.
Starting from an input, we ran the loop of the ItrSA model for $128$ steps to obtain $x_{t=0}$. We then repeated the run from a perturbed input $x_{t=0}+\epsilon$, where $\epsilon\sim N(0,10^{-3})$ is added at $t=0$. From each trajectory, we sampled $300$ equally spaced points between $t=0$ and $t=10000$. The results are shown in Figure˜S.5.
Figure˜S.5 plots the L1 distance between the two trajectories. In Figures˜S.5 and S.5, we visualize a single coordinate of five trajectories. Figure˜S.5 shows the complete trajectories. Since we observed that the trajectories rapidly oscillate between two clusters and the visualization was subtle, Figure˜S.5 provides a zoomed-in view of one cluster (values >0.1>0.1).
Overall, the distance between the original and perturbed trajectories increases exponentially over time, demonstrating chaotic behavior in the system.
(a) L1 distance between trajectories
(b) All sampled trajectories
(c) Trajectories forming a single cluster
Figure S.5: Sensitivity to initial conditions investigated using the ItrSA model.
B.5.5 Jacobian-based interpretation of pseudo-energy
We provide additional insights into the interpretation of pseudo-energy via the Jacobian discussed in Section˜6.3. Using Lemma A.3 from Noci et al. [2022], the Jacobian of SA is expressed as
@@ B.5.5 Jacobian-based interpretation of pseudo-energy
Figure˜S.6 empirically confirmed that both the contribution index and the pseudo-energy behave similarly even under these approximations. PV indicates the contribution index using the approximation (117), and PV ($T=0$) uses (118). The figure shows that the contribution index can be effectively explained using only the attention matrix at the initialization of inference. The blue curve shows the approximated pseudo-energy (119) and works similarly to the original one. Thus, our Jacobian-based analysis interprets the pseudo-energy as quantifying exploration in eigenspaces corresponding to large eigenvalues of the attention matrix determined by the initial inference states.
@@
B.5.6 CIFAR-10
For the experiments on the CIFAR-10 dataset [Krizhevsky et al., ], we used the same architecture and setup as in the Sudoku experiments. We used the Adam optimizer and tuned the learning rate across $\{1\times 10^{-6},5\times 10^{-6},\dotsc,1\times 10^{-3}\}$ based on the test accuracy at the iteration $T=16$. We trained models for 200 epochs and set the batch size $128$. We used $N=4$ as the dimension of oscillators of AKOrN.
Figure˜S.7 shows the Lyapunov exponent on the CIFAR-10 dataset. This result is in the same trend with that on the Sudoku dataset.
@@ C.1 Extended related work
The Transformer architecture has been a focus of efforts to provide theoretical grounding. Geshkovski et al. [2025, 2023, 2024] formulated recurrent SA dynamics as interactions among tokens (“particles”), enabling theoretical analysis of phenomena such as meta-stable clustering and rank-one collapse. Their continuous-time dynamics monotonically decrease an energy (Lyapunov) function given by a summation over exponential functions, commonly requiring constraints such as single-head attention, hyperspherical token states, and symmetric weights. Karagodin et al. [2024] extended this framework to the case of causal attention masking. Bruno et al. [2025] succeeded in mathematically characterizing the meta-stable clustering as a Wasserstein gradient flow of mean-field token dynamics, with the energy serving as its potential function, although they replaced the softmax function with an unnormalized exponential function and restricted their analysis to identity weight matrices. Yang et al. [2022] considered an exponential energy function similar to that of Geshkovski et al. [2025], describing the Transformer as performing alternating majorization-minimization updates on distinct energy functions. Their approach also accommodates discrete state updates and MLP layers, although it entails complex conditions, including constraints on step sizes and proximity to fixed points. Ramsauer et al. [2021] formalized the cross-attention mechanism as modern Hopfield networks, Hoover et al. [2023], Hu et al. [2025] further developed energy functions for Transformers including self-attentions. We do not address approaches based on Hopfield networks in this work, as they require architectural modifications, such as adding auxiliary signal paths that are absent in standard Transformers, which are beyond our scope.